# Cooperative Bistatic ISAC Systems for Low-Altitude Economy

### v1

@@ Abstract
Index Terms:
Cooperative integrated sensing and communication (ISAC), orthogonal frequency division multiplexing (OFDM), low-altitude economy (LAE), bistatic sensing.
@@ I Introduction
The rapid evolution toward sixth-generation (6G) and NextG wireless networks has ushered in a transformative vision where communication systems transcend traditional data transmission to embrace advanced sensing capabilities [1]. Integrated sensing and communication (ISAC) is central to realizing this vision, unifying sensing and communication within shared infrastructure. This convergence promises significant gains for optimized spectral efficiency, reduced hardware costs, and novel application domains [2]. By exploiting the inherent synergy between sensing and communication, ISAC not only enhances network adaptability but also paves the way for intelligent environments where real-time situational awareness coexists with ultra-reliable connectivity. ISAC is particularly transformative for the low-altitude economy (LAE), where its capabilities align directly with critical operational needs [3]. LAE encompasses applications spanning logistics, agriculture, and emergency services, which rely on unmanned aerial vehicles (UAVs) operating in complex airspace, demanding robust detection, localization, and tracking capabilities alongside seamless connectivity. Addressing these requirements necessitates ISAC protocols tailored to LAE’s unique challenges, including high-accuracy multi-target sensing and compatibility with the existing cellular infrastructure.
@@
Similar to radar systems, ISAC implementations typically adopt two basic operational modes: monostatic (a.k.a. active sensing) [4, 5] and bistatic (a.k.a. passive sensing) [6]. In monostatic systems, a single base station (BS) simultaneously transmits and receives sensing signals, enabling independent operation. However, this approach requires full-duplex transceivers with self-interference mitigation capability, which is not supported by the existing commercial BSs, necessitating costly hardware upgrades. Conversely, bistatic configurations separate the transmitter and receiver, enabling sensing through reflected signals. This sensing pattern broadens coverage and avoid costly full-duplex hardware, facilitating integration with the existing systems [7, 8], while introducing synchronization requirements between BSs. In addition to these challenges, conventional configurations-whether monostatic or bistatic-also have inherent limitations: sparse spatial coverage due to limited transmission power and environmental obstacles, speed estimation restricted to the radial component, and susceptibility to measurement outliers.
@@
Cooperative ISAC architectures, inspired by multistatic radar principles, are expected to overcome these limitations [9, 10]. By leveraging distributed BSs to collaboratively transmit and receive sensing signals, cooperative ISAC systems amplify coverage, enhance measurement diversity, and improve robustness against environmental dynamics. In cooperative ISAC networks, multiple BSs collaboratively transmit and receive signals, generating rich spatial-temporal measurements. Sensing parameters such as time delays, Doppler shifts, and angles are extracted locally at BSs or centrally at a central processing unit (CPU). These parameters are then fused at the CPU for high-accuracy target localization and true velocity reconstruction. In our previous work [11], a cell-free-based cooperative bistatic ISAC framework was proposed. It revealed that such cooperation can be achieved through protocol-level innovations, such as staggered uplink/downlink scheduling across BSs, which repurposed existing orthogonal frequency-division multiplexing (OFDM) communication waveforms and infrastructure for cooperative sensing without hardware updates. This alignment with the standardized 5G New Radio (NR) frame structure demonstrated how cooperative ISAC combines theoretical advancements with practical deployment capabilities, which is a key consideration for 6G standardization.
@@
Prior research has established foundational frameworks for cooperative ISAC-enabled sensing schemes. For example, [11] and [12] addressed single-antenna BS configurations, employing bistatic and monostatic sensing modes through two-dimensional (2D) fast Fourier transform (2D-FFT) and compressed sensing algorithms for target localization. Subsequent studies extended these concepts to multi-antenna architectures. For the scenario with multi-antenna BSs, Lu et al. [13] investigated a multi-BS cooperative system using multiple signal classification (MUSIC) algorithms for parameter estimation, complemented by a lattice point search method for data fusion. This progression continued with Wei et al. [14], who developed a symbol-level data fusion scheme integrating phase features in the demodulation symbols with distance and radial velocity estimations in cooperative monostatic ISAC systems, later adapted by the same research group [15] for bistatic configurations with enhanced position and velocity estimation capabilities based on angle-of-arrival (AoA) and angle-of-departure (AoD) extractions. In addition, feasible topology and frame structure designs were proposed by Han et al. [16], which achieved cooperative bistatic sensing in the cellular networks at a low cost. A hybrid monostatic-bistatic architecture was proposed by Jiang et al. [17], where dual BSs jointly process echo signals from both BSs for improved target localization. Despite these advances, most studies focused on simplified scenarios, such as single-antenna configurations [12, 11], single-target sensing [13, 14, 15, 16], and 2D spatial estimation [12, 11, 14, 15, 17, 16], which failed to reflect the complexity of practical deployments like LAE.. Emerging solutions attempted to address these constraints through advanced mathematical tools. Tang et al. [18] introduced tensor decomposition techniques to cooperative monostatic systems, enabling low-complexity parameter estimation of range, angle, and velocity with inherent multi-dimensional parameter pairing. This approach has already been validated in channel estimation [19, 20], showing promise for the ISAC-enabled sensing scheme designs [21].
@@
Given the extensive coverage of mobile networks, developing sensing schemes compatible with existing infrastructure is crucial for practical ISAC deployment and application proliferation. However, two key limitations hinder practical deployment: First, as mentioned above, despite the prevalence of massive MIMO BSs, existing studies predominantly addressed simplified single-antenna [12, 11] or single-target scenarios [13, 14, 15, 16], failing to leverage modern network capabilities. Second, bistatic architectures [11, 15, 16] have inherent advantages over monostatic approaches [12, 14, 13, 18] in hardware compatibility by reusing the existing communication equipment, while also generating richer sensing data. For instance, with 5 cooperative BSs configured as 2 transmitters and 3 receivers, bistatic sensing schemes can generate $2\times 3$ measurement links, surpassing the 5 self-transceiving links in monostatic schemes. Meanwhile, the reduction in the number of transmitters also decreases spectrum consumption and inter-BS interference. Nevertheless, fusing abundant bistatic measurements to enhance sensing accuracy remains challenging. While [11] proposed a solution for single-antenna scenarios, its high complexity limits its scalability in environments with a large number of BSs or targets. Although the data fusion framework in [18] serves as a viable paradigm for monostatic systems, it cannot be directly applied to bistatic architectures due to the inherent challenges in bistatic sensing, such as the elimination of symbol timing offset (STO) and carrier frequency offset (CFO), as well as issues related to three-dimensional (3D) position estimation.
@@
    We develop a low-complexity parameter estimation algorithm using ESPRIT-inspired CANDECOMP/PARAFAC (CP) tensor decomposition, which efficiently extracts range, velocity, and angle parameters of the targets by leveraging the Vandermonde structure of delay-related factor matrices.
@@ II-A Signal Transmission Model
As illustrated in Fig. 1, we consider a MIMO-OFDM network-based cooperative ISAC system capable of detecting, locating, and measuring velocities of low-altitude targets, thereby supporting the low-altitude economy applications. This system employs a bistatic sensing architecture where $N_{\mathrm{t}}$ transmit BSs (tBSs) simultaneously emit sensing signals, while $N_{\mathrm{r}}$ receive BSs (rBSs) collaboratively estimate the parameters of $K$ targets through processing the reflected echoes.111 While ISAC systems can dynamically reconfigure tBS/rBS assignments based on specific sensing requirements, this work focuses exclusively on developing sensing algorithms, with BS selection strategies proposed in [22]. Operating in the time division duplexing (TDD) mode, all BSs are connected with a CPU through fronthaul links. The CPU sends timing information to the BSs via the fronthaul links for synchronization, performs data fusion of the rBS-derived sensing information to enhance the sensing performance, and adjusts the slot format [23] of each BS to facilitate the sensing signal exchange. A detailed example of the slot format configuration is presented in Fig. 2. Similar to the cooperative ISAC framework established in our previous work [11], to achieve bistatic sensing in mobile communication systems, the rBSs are configured with slot formats different from the tBSs when performing sensing tasks. During the $N$-symbol sensing task, each tBS transmits pre-determined sensing reference signals on non-overlapping bandwidth parts (BWPs) to eliminate the inter-BS interference, enabling unambiguous signal separation and identification at all rBSs. The estimated parameters from these signals are then uploaded to the CPU through the fronthaul links.
@@
Within the common 5G telecommunication infrastructure, every BS implements a MIMO-OFDM air interface employing uniform planar arrays (UPAs) for 3D beamforming. Specifically, the UPAs are installed with an elevation angle of 0° to facilitate both the sensing of low-altitude targets and the communication service for terrestrial users. To achieve cost-efficient massive antenna deployment, we consider an architecture that all BSs adopt the full-connected hybrid beamforming structure, which reduce the number of radio frequency (RF) chains. We assume that each BS comprises $R$ RF chains driving $L=N_{\mathrm{h}}\times N_{\mathrm{v}}>R$ antenna elements, where $N_{\mathrm{v}}$ and $N_{\mathrm{h}}$ denote the numbers of antenna elements in the vertical and horizontal directions, respectively.
In the considered sensing framework, all rBSs independently execute identical processing algorithms on signals transmitted from individual tBSs. In the following, we consider the signal received at rBS $n_{\mathrm{r}}$ as an illustrative case for presenting the transmission model and the proposed parameter extraction scheme. For notational compactness, we omit the subscripts $n_{\mathrm{t}}$ and $n_{\mathrm{r}}$ throughout the derivations in the following. Let $s_{l,m,n}$ denote the modulated frequency-domain sensing symbol allocated in the $m$th subcarrier and the $n$th OFDM symbol at the $l$th transmit antenna, which satisfies $\left|s_{l,m,n}\right|^{2}=1$. Then, the signal transmitted by tBS $n_{\mathrm{t}}$ can be represented as
$s_{l}\left(t\right)=\sum_{m=1}^{M}{\sum_{n=1}^{N}{s_{l,m,n}}e^{j2\pi m\Delta ft}}\cdot r\left(t-nT_{\mathrm{s}}\right),\,l=1,...,L,$ (1)
where $M$ is the number of subcarriers, $T_{\mathrm{s}}$ denotes the OFDM symbol period (including the cyclic prefix), $r(t)$ denotes the pulse-shaping filter function, and $\Delta f$ denotes the subcarrier spacing (SCS). Following matched filtering, cyclic prefix removing, and inverse fast Fourier transform (IFFT), the frequency-domain signal vector received at rBS $n_{\mathrm{r}}$ during the $n$th sensing symbol is expressed as
@@
where $\mathbf{H}_{m,n}\in\mathbb{C}^{L\times L}$ denotes the discrete frequency-domain channel (its formulation will be derived in the subsequent subsection), the hybrid precoding matrix $\mathbf{F}\triangleq\mathbf{F}_{\mathrm{a}}\mathbf{F}_{\mathrm{d}}$ combines an analog phase-shifting network $\mathbf{F}_{\mathrm{a}}\in\mathbb{C}^{L\times R}$ with a digital precoder $\mathbf{F}_{\mathrm{d}}\in\mathbb{C}^{R\times R}$, while the hybrid combining matrix $\mathbf{Q}=\mathbf{Q}_{\mathrm{a}}\mathbf{Q}_{\mathrm{d}}$ integrates an analog combining matrix $\mathbf{Q}_{\mathrm{a}}\in\mathbb{C}^{L\times R}$ and a digital combining matrix $\mathbf{Q}_{\mathrm{d}}\in\mathbb{C}^{R\times R}$. The transmitted symbol vector $\mathbf{s}_{m,n}=[s_{1,m,n},\cdots,s_{l,m,n},\cdots,s_{L,m,n}]^{\mathrm{T}}\in\mathbb{C}^{L\times 1}$ contains the sensing symbols from tBS $n_{\mathrm{t}}$, and $\mathbf{n}_{m,n}\in\mathbb{C}^{L\times 1}$ represents the additive white Gaussian noise vector. Given the known symbol vector $\mathbf{s}_{m,n}$, the rBSs can eliminate the impact of it by multiplying the received signal vector by $\mathbf{s}_{m,n}^{*}$, i.e.,
$\hat{\mathbf{y}}_{m,n}=\mathbf{s}_{m,n}^{*}*\mathbf{y}_{m,n}=\mathbf{Q}^{\mathrm{H}}\mathbf{H}_{m,n}\mathbf{F1}+\hat{\mathbf{n}}_{m,n},$ (3)
@@
II-B Sensing Channel Model

### v2

@@ I Introduction
The rapid evolution toward sixth-generation (6G) and NextG wireless networks has ushered in a transformative vision where communication systems transcend traditional data transmission to embrace advanced sensing capabilities [1]. Integrated sensing and communication (ISAC) is central to realizing this vision, unifying sensing and communication within shared infrastructure. This convergence promises significant gains for optimized spectral efficiency, reduced hardware costs, and novel application domains [2]. By exploiting the inherent synergy between sensing and communication, ISAC not only enhances network adaptability but also paves the way for intelligent environments where real-time situational awareness coexists with ultra-reliable connectivity. ISAC is particularly transformative for the low-altitude economy (LAE) and its supporting low-altitude wireless networks (LAWNs), where its capabilities align directly with critical operational needs [3]. LAE encompasses applications spanning logistics, agriculture, and emergency services, which rely on unmanned aerial vehicles (UAVs) operating in complex airspace, demanding robust detection, localization, and tracking capabilities alongside seamless connectivity. Addressing these requirements necessitates ISAC protocols tailored to LAE’s unique challenges, including high-accuracy multi-target sensing and compatibility with the existing cellular infrastructure.
@@
Similar to radar systems, ISAC implementations typically adopt two basic operational modes: monostatic (a.k.a. active sensing) [4, 5] and bistatic (a.k.a. passive sensing) [6]. In monostatic systems, a single base station (BS) simultaneously transmits and receives sensing signals, enabling independent operation. However, this approach requires full-duplex transceivers with self-interference mitigation capability, which is not supported by the existing commercial BSs, necessitating costly hardware upgrades. Conversely, bistatic configurations separate the transmitter and receiver. This sensing pattern broadens coverage and avoids costly full-duplex hardware, facilitating integration with the existing systems [7, 8], while introducing synchronization requirements between BSs. In addition to these challenges, conventional configurations-whether monostatic or bistatic-also have inherent limitations: sparse spatial coverage due to limited transmission power and environmental obstacles, speed estimation restricted to the radial component, and susceptibility to measurement outliers.
@@
Cooperative ISAC architectures, inspired by multistatic radar principles, are expected to overcome these limitations [9, 10]. By leveraging distributed BSs to collaboratively transmit and receive sensing signals, cooperative ISAC systems amplify coverage, enhance measurement diversity, and improve robustness against environmental dynamics. In cooperative ISAC networks, multiple BSs can generate rich spatial-temporal measurements. Sensing parameters such as time delays, Doppler shifts, and angles are extracted locally at BSs or centrally at a central processing unit (CPU). These parameters are then fused at the CPU for high-accuracy target localization and true velocity reconstruction. In our previous work [11], a cell-free-based cooperative bistatic ISAC framework was proposed. It revealed that such cooperation can be achieved through protocol-level innovations, such as staggered uplink/downlink scheduling across BSs, which repurposed existing orthogonal frequency-division multiplexing (OFDM) communication waveforms and infrastructure for cooperative sensing without hardware updates. This alignment with the standardized 5G New Radio (NR) frame structure demonstrated how cooperative ISAC combines theoretical advancements with practical deployment capabilities, which is a key consideration for 6G standardization.
@@
Prior research has established foundational frameworks for cooperative ISAC-enabled sensing schemes. For example, [11] and [12] addressed single-antenna BS configurations, employing bistatic and monostatic sensing modes through two-dimensional (2D) fast Fourier transform (2D-FFT) and compressed sensing algorithms for target localization. Subsequent studies extended these concepts to multi-antenna architectures. For the scenario with multi-antenna BSs, Lu et al. [13] investigated a multi-BS cooperative system using multiple signal classification (MUSIC) algorithms for parameter estimation, complemented by a lattice point search method for data fusion. This progression continued with Wei et al. [14], who developed a symbol-level data fusion scheme integrating phase features in the demodulation symbols with distance and radial velocity estimations in cooperative monostatic ISAC systems, later adapted by the same research group [15] for bistatic configurations with enhanced position and velocity estimation capabilities based on angle-of-arrival (AoA) and angle-of-departure (AoD) extractions. In addition, feasible topology and frame structure designs were proposed by Han et al. [16], which achieved cooperative bistatic sensing in the cellular networks at a low cost. A hybrid monostatic-bistatic architecture was proposed by Jiang et al. [17], where dual BSs jointly process echo signals from both BSs for improved target localization. Emerging solutions attempted to address constraints in simplified scenarios through advanced mathematical tools. Our previous work [18] introduced tensor decomposition techniques to cooperative monostatic systems, enabling low-complexity parameter estimation of range, angle, and velocity with inherent multi-dimensional parameter pairing. This approach has already been used in channel estimation [19, 20], showing promise for the ISAC-enabled sensing scheme designs [21].
@@
Given the extensive coverage of mobile networks, developing sensing schemes compatible with existing infrastructure is key for practical ISAC deployment and application spread. However, despite the prevalence of massive MIMO BSs, existing studies mainly addressed simplified scenarios, such as single-antenna configurations [12, 11], single-target sensing [13, 14, 15, 16], and 2D spatial estimation [12, 11, 14, 15, 17, 16], failing to fully leverage network capabilities or reflect the complexity of practical deployments like LAE. In this context, bistatic architectures [11, 15, 16] have inherent advantages over monostatic approaches [12, 14, 13, 18] in hardware compatibility by reusing the existing communication equipment, while also generating richer sensing data. For instance, with 5 cooperative BSs configured as 2 transmitters and 3 receivers, bistatic schemes can generate $2\times 3$ measurement links, exceeding the 5 self-transceiving links in monostatic schemes. Meanwhile, fewer transmitters reduce spectrum resource consumption and inter-BS interference. Nevertheless, fusing abundant bistatic measurements to enhance sensing accuracy remains challenging. While [11] proposed a solution for single-antenna scenarios, the high complexity limits its scalability in environments with a large number of BSs or targets. Although the data fusion framework in [18] provided a viable paradigm for monostatic systems, it cannot be directly applied to bistatic architectures due to inherent challenges in bistatic sensing. For example, symbol timing offset (STO) and carrier frequency offset (CFO) need to be effectively suppressed, otherwise serious range and velocity estimation errors will occur. In addition, due to the geographical separation between transmitters and receivers in bistatic sensing, the angle of departure (AOD) and angle of arrival (AOA) related to a target are different, and the distance between a target and a BS cannot be directly estimated, which increases the difficulty of parameter estimation and data fusion.
@@
    We develop a low-complexity parameter estimation algorithm using Estimation of Signal Parameters via Rotational Invariance Techniques (ESPRIT)-inspired CANDECOMP/PARAFAC (CP) tensor decomposition, which efficiently extracts range, velocity, and angle parameters of the targets by leveraging the Vandermonde structure of delay-related factor matrices.
@@ II-A Signal Transmission Model
As illustrated in Fig. 1, we consider a MIMO-OFDM network-based cooperative ISAC system capable of detecting, locating, and measuring velocities of low-altitude targets, thereby supporting the low-altitude economy applications. This system employs a bistatic sensing architecture, where $N_{\mathrm{t}}$ transmit BSs (tBSs) simultaneously emit sensing signals, while $N_{\mathrm{r}}$ receive BSs (rBSs) collaboratively estimate the parameters of $K$ targets through processing the reflected echoes. [^1] While ISAC systems can dynamically reconfigure tBS/rBS assignments based on specific sensing requirements, this work focuses exclusively on developing sensing algorithms, with BS selection strategies proposed in [22]. Operating in the time division duplexing (TDD) mode, all BSs are connected with a CPU through fronthaul links. The CPU sends timing information to the BSs via the fronthaul links for synchronization, performs data fusion of the rBS-derived sensing information to enhance the sensing performance, and adjusts the slot format [23] of each BS to facilitate the sensing signal exchange. A detailed example of the slot format configuration is presented in Fig. 2. Similar to the cooperative ISAC framework established in our previous work [11], to achieve bistatic sensing in mobile communication systems, the rBSs are configured with slot formats different from the tBSs when performing sensing tasks. During the $N$-symbol sensing task, each tBS transmits pre-determined sensing reference signals on non-overlapping bandwidth parts (BWPs) to eliminate the inter-BS interference. [^2] To minimize modifications to existing transmission protocols, dedicated symbol resources are allocated for sensing by suspending communication services. Although this incurs throughput loss, the impact is negligible since sensing tasks require only a few symbols and occur at relatively low refresh rates (e.g., once every few seconds or a few times per second). The estimated parameters from these signals are then uploaded to the CPU through fronthaul links.
@@
Within the common 5G telecommunication infrastructure, every BS implements a MIMO-OFDM air interface employing uniform planar arrays (UPAs) for 3D beamforming [24]. Specifically, the UPAs are installed with an elevation angle of 0° to facilitate both the sensing of low-altitude targets and the communication service for terrestrial users. To achieve cost-efficient massive antenna deployment, we consider an architecture that all BSs adopt the fully-connected hybrid beamforming structure, which reduce the number of radio frequency (RF) chains. We assume that each BS comprises $R$ RF chains driving $L=N_{\mathrm{h}}\times N_{\mathrm{v}}>R$ antenna elements, where $N_{\mathrm{v}}$ and $N_{\mathrm{h}}$ denote the numbers of antenna elements in the vertical and horizontal directions, respectively.
Note that each rBS can effectively separate and identify the transmitted signals from each tBS by applying bandpass filtering to each BWP. In addition, all rBSs in the considered sensing framework independently execute identical processing algorithms on signals transmitted from individual tBSs. Therefore, we consider the signal transmitted from tBS $n_{\mathrm{t}}$ to rBS $n_{\mathrm{r}}$ as an illustrative case for presenting the transmission model and the proposed parameter extraction scheme. For notational compactness, we omit the subscripts $n_{\mathrm{t}}$ and $n_{\mathrm{r}}$ throughout the derivations in the following. Let $s_{r,m,n}$ denote the modulated frequency-domain sensing symbol allocated in the $m$th subcarrier of the $n$th OFDM symbol input to the $r$th RF chain, which satisfies $\left|s_{r,m,n}\right|^{2}=1$. Then, the baseband time-domain OFDM signal transmitted by tBS $n_{\mathrm{t}}$ can be represented as
$s_{r}\left(t\right)=\sum_{m=1}^{M}{\sum_{n=1}^{N}{s_{r,m,n}}e^{j2\pi m\Delta ft}}\cdot r\left(t-nT_{\mathrm{s}}\right),\,l=1,...,L,$ (1)
where $M$ is the number of subcarriers, $T_{\mathrm{s}}$ denotes the OFDM symbol period (including the cyclic prefix), $r(t)$ denotes the pulse-shaping filter function, and $\Delta f$ denotes the subcarrier spacing (SCS). Following matched filtering, cyclic prefix removing, and inverse fast Fourier transform (IFFT), the frequency-domain signal vector received at the rBS during the $n$th sensing symbol is expressed as
@@
where $\mathbf{H}_{m,n}\in\mathbb{C}^{L\times L}$ denotes the discrete frequency-domain channel (its formulation will be derived in the subsequent subsection), the hybrid precoding matrix $\mathbf{F}\triangleq\mathbf{F}_{\mathrm{a}}\mathbf{F}_{\mathrm{d}}$ combines an analog phase-shifting network $\mathbf{F}_{\mathrm{a}}\in\mathbb{C}^{L\times R}$ with a digital precoder $\mathbf{F}_{\mathrm{d}}\in\mathbb{C}^{R\times R}$, while the hybrid combining matrix $\mathbf{Q}=\mathbf{Q}_{\mathrm{a}}\mathbf{Q}_{\mathrm{d}}$ integrates an analog combining matrix $\mathbf{Q}_{\mathrm{a}}\in\mathbb{C}^{L\times R}$ and a digital combining matrix $\mathbf{Q}_{\mathrm{d}}\in\mathbb{C}^{R\times R}$. The transmitted symbol vector $\mathbf{s}_{m,n}=[s_{1,m,n},\cdots,s_{r,m,n},\cdots,s_{R,m,n}]^{\mathrm{T}}\in\mathbb{C}^{L\times 1}$ contains the sensing symbols from tBS $n_{\mathrm{t}}$, and $\mathbf{n}_{m,n}\in\mathbb{C}^{L\times 1}$ represents the additive white Gaussian noise vector. Given the known symbol vector $\mathbf{s}_{m,n}$, the rBSs can eliminate the impact of it by multiplying the received signal vector by $\mathbf{s}_{m,n}^{*}$, i.e.,
$\hat{\mathbf{y}}_{m,n}=\mathbf{y}_{m,n}*\mathbf{s}_{m,n}^{*}=\mathbf{Q}^{\mathrm{H}}\mathbf{H}_{m,n}\mathbf{F1}+\hat{\mathbf{n}}_{m,n},$ (3)
@@
II-B Beanforming for Low-altitude Targets
In practical sensing scenarios, the positions of low-altitude targets are typically unknown a priori. An Intuitive method to enhance the echo signal-to-noise ratio (SNR) and improve sensing accuracy is to perform beam scanning to obtain coarse directional information of the targets, followed by steering beams toward these directions [18]. However, due to the narrow beamwidth, exhaustive beam scanning using large antenna arrays incurs a substantial time overhead. Conversely, random beamforming tends to spread transmit energy excessively across space, as illustrated in Fig. 3(a), leading to weak target echoes and strong ground clutter. To balance beamforming gain and clutter suppression, an appropriate design of the beamforming/combining matrices $\mathbf{F}$ and $\mathbf{Q}$ is essential. Based on the above discussions, an ideal beamforming scheme should concentrate radiated energy within the sensing region of interest. Specifically, let $\mathcal{R}$ denote the angular region defined by elevation range $\theta_{\min},\theta_{\max}]$ and azimuth range $\phi_{\min},\phi_{\max}]$. The beamforming gain $\mathrm{Gain}(\theta,\phi)$ should satisfy:
[…]
where $\varepsilon_{0}\ll\mathrm{Gain}_{0}$ to suppress ground clutter. That is, the gain should be approximately uniform and high within $\mathcal{R}$, and minimized outside it. Although precoding design is a well-studied topic and not the focus of this work, we provide an exemplary design for the precoding matrix $\mathbf{F}$ to illustrate that such beam patterns are feasible in practice. A similar approach can be applied to construct the combining matrix $\mathbf{Q}$.
Consider a sensing region defined by elevation range from 40° to 90° and azimuth range from 40° to 140°. With $R=64$ RF chains, we uniformly sample this region into 64 angle pairs, using 8 elevation samples $\left[\theta_{1}^{\mathrm{samp}},\dots,\theta_{8}^{\mathrm{samp}}\right]=\left[40,47.14,\dots,90\right]$ and 8 azimuth samples $\left[\phi_{1}^{\mathrm{samp}},\dots,\phi_{8}^{\mathrm{samp}}\right]=\left[40,54.29,\dots,140\right]$. Let $\mathbf{F}_{\mathrm{d}}=\mathbf{I}$ and construct $\mathbf{F}_{\mathrm{a}}$ such that its $\left(p_{\mathrm{s}}+8(q_{\mathrm{s}}-1)\right)$-th column is given by the transmit steering vector $\mathbf{a}_{\mathrm{t}}\left(\theta_{p_{\mathrm{s}}}^{\mathrm{samp}},\phi_{q_{\mathrm{s}}}^{\mathrm{samp}}\right)$, whose definition will be given in the next subsection. As shown in Fig. 3(b), this design effectively suppresses energy leakage toward the ground.
It is worth emphasizing that the above example is neither the only feasible design nor necessarily optimal. Nevertheless, it demonstrates that by properly configuring beamforming matrices $\mathbf{F}$ and $\mathbf{Q}$ at the tBSs and rBSs, ground clutter can be significantly suppressed. Therefore, in the subsequent derivations of this paper, we ignore the ground clutter component in the channel $\mathbf{H}_{m,n}$.
(a) Random
(b) Designed
Figure 3: Radiation patterns achieved by random beamforming and the exemplary design for low-altitude target sensing.
II-C Sensing Channel Model